Is it true or false Given 1 L = 1 dm3 and 1 mL = 1 cm3
1
Expert's answer
2020-08-30T02:22:36-0400

It's true.

"1 ~m^3" is equal to 1000 L, so "1~dm^3=10^{-3}~m^3=10^{-3}*1000~L=1~L"

"1~ml=10^{-3}~L=10^{-3}~dm^3=10^{-3}*10^{-3}~m^3=10^{-6}~m^3"

"1~m = 10^2~cm, so~1~m^3=10^6~cm^3"

Finally, "1~ml=10^{-6}~m^3=10^{-6}*10^6~cm^3=1~cm^3"